# A Survey on State-of-the-art Deep Learning Applications and Challenges

### v5

@@ 4.1.3 Image Segmentation
Image segmentation is another important task in which deep learning has a significant impact. One of the earliest deep learning models for image segmentation is the fully convolutional network [110]. A fully convolutional network consists of only convolutional layers which accepts an input of an arbitrary size and produce the predicted segmentation map of the same size. The authors adopt the AlexNet, VGG16 and GoogleNet, replace their fully connected layers with convolutional layers and append a 1×1 convolutional layer, followed by bilinear up-sampling to match the size of the input. The model was considered a significant milestone in image segmentation, demonstrating the feasibility of deep learning for semantic segmentation trained in end-to-end manner. Deconvolution network is another popular deep learning model for semantic segmentation [111]. The model architecture consists of two parts: encoder and decoder. The encoder takes an input image and uses the convolutional layers to generate the feature maps. The feature maps are fed to the decoder composed of un-sampling and deconvolutional layers to predict the segmentation map. SegNet is another encoder-decoder model for semantic segmentation [112]. The encoder is a sequence of convolutional (with ReLU) and max-pooling blocks which is analogous to a convolutional neural network. The decoder is composed of up-sampling layers which up-samples the inputs using the memorized pooled indices generated in the encoder phase, and convolutional layers without non-linearity. The encoder progressively reduces the resolution of the input data while extracting abstract features through a series of convolutional and pooling layers. This process causes the loss of fine-grained information, degrading the overall performance of segmentation. LinkNet mitigates this limitation by passing the feature maps at several stages generated by the encoder to the decoder, hence reducing information loss [113]. The model architecture of LinkNet is similar to SegNet, but utilizes ResNet as the encoder.
@@
While Faster R-CNN is a significant approach in object detection task, it has been extended to perform instance segmentation task. One such extension is Mask R-CNN which is based on Faster R-CNN, introduces an additional branch for predicting the segmentation mask [114]. Similar to Faster R-CNN, Mask R-CNN utilizes the RPN to generate region proposals and then the region of interest alignment is applied to extract more accurate features from the proposed regions. Mask R-CNN does not leverage the multiscale feature representation which may degrade the overall performance of segmentation. To overcome this limitation, Path Aggregation Network (PANet) incorporates the FPN and introduces a bottom-up pathway to facilitate the propagation of the low-level information [115]. The pathway takes the feature maps of the previous stage as input and performs 3×3 convolution with stride 2 to reduce the spatial size of the feature maps. The generated feature maps are then fused with the feature maps from the FPN through the lateral connection. The model adopts the three branches as in Mask R-CNN. MaskLab is an instance segmentation model based on Faster R-CNN, consisting of object detection, segmentation and instance (object) center direction prediction branches [116]. The direction prediction provides useful information to distinguish instances of the same semantic label, allowing the model to further refine the instance segmentation results.
@@
The convolutional layers use local receptive fields to process input data which can be effective for exploiting spatial patterns and hierarchical features but may find it difficult to capture global relationships across the entire image. The ViT has been leveraged to mitigate this issue in semantic segmentation [119]. Specifically, the input image is divided into patches and treated as input to the transformer to capture the global relationship between the patches, significantly improving the prediction of the segmentation map. Global context ViT aims to address the lack of ViT’s ability to leverage local spatial features [120]. As shown in Figure 20, the transformer consists of local and global self-attention modules. The role of global self-attention is to capture the global contextual information from different image regions while the short-range information is captured by the local self-attention. Multiscale feature representation is crucial for accurate semantic segmentation. However, the transformer often combines the features without considering their appropriate (optimal) scales, thus affecting the segmentation accuracy. Transformer scale gate is a module proposed to address the issue of selecting an appropriate scale based on the correlation between patch-query pairs [121]. The transformer takes attention (correlation) maps as input and calculates the weights of the multi-scale features for each image patch, allowing the model to adaptively choose the optimal scale for each patch.
@@ 4.1.4 Image Generation
Image generation refers to the process of creating images based on input texts. Generally, the task can be divided into three stages. The first stage is extracting features from the input text, followed by generating the image and finally controlling the image generation process to ensure the output meets specific criteria and constraints. This section focuses on the progress made in the development of deep learning models of the second stage (image generation) since it directly impacts the quality of the generated images. Variational autoencoder is one of the earliest deep learning models that is capable of generating images [67]. Variational autoencoder learns to generate data by capturing the underlying (Gaussian) distribution of the training data. During the generation process, the distribution parameters are sampled and passed to the decoder to generate the output image. Although the generated images are blurry and unsatisfactory, it has shown a lot of potential in image generation tasks. The introduction of GAN significantly improved the quality of generated images. GAN consists of two connected neural networks, a generator and a discriminator that are trained simultaneously in a competitive manner [68]. The generator learns to generate realistic images to fool the discriminator, while the discriminator learns to distinguish between fake and real images. The generated images are less blurry and more realistic. Several enhanced models have been proposed to improve its usability and overall performance such as CGAN [69] which allows us to tell what image to be generated, and the deep convolutional GAN (DCGAN) [122] which provides a more stable structure for image generation. DCGAN is the basis of many subsequent improvements in GANs.
@@
Although multistage generators improve image generation performance by leveraging multiscale representation, the generated images may contain fuzzy shapes with coarse features. DF-GAN replaces the multistage generators with a single-stage deep generator featuring residual connections and trained with hinge loss [130]. Furthermore, DF-GAN introduces a regularization strategy on the discriminator that applies a gradient penalty on real images with matching text, allowing the model to generate more text-matching images. DMF-GAN an improved DF-GAN, incorporates three novel components designed to leverage semantic coherence between the input text and the generated image [131]. The first component is the recurrent semantic fusion module, which models long range dependencies between the fusion blocks. The second component is the multi-head attention module which is placed towards the end of the generator to leverage the word features, forcing the generator to generate images conditioned on the relevant words. The last component is the word-level discriminator which provides fine-grained feedback to the generator, facilitating the learning process and improving the overall quality of the generated images. Figure 22 shows the architecture of DMF-GAN. The process of image generation involves feeding a noise vector to the generator at the very beginning of the network. However, as the generator goes deeper, the noise effect may be diminished, affecting the diversity of the image generation results. To mitigate this issue, DE-GAN incorporates a dual injection module into the single-stage generator [132]. The dual injection module consists of two text fusion layers followed by a noise broadcast operation. The text fusion layer takes the sentence embedding and fuses it with the input feature map using the fully-connected layer. Then noise is injected into the output feature map to retain the randomness in the generation process, improving diversity and generalization of the model.
@@ 4.2 Natural Language Processing
Natural language processing (NLP) refers to the field of artificial intelligence that concerns with enabling computers to process, analyze and interpret human languages to extract useful information. Some of the common tasks in NLP are machine translation, text classification and text generation. Deep learning has been widely applied to solve real-world NLP problems. This section presents the recent advancements in deep learning models that have been designed for NLP over the past few years.
@@ 4.2.1 Text Classification
Text classification known as text categorization, is a task that involves assigning predefined categories or labels to a piece of text based on its content. The task is commonly used in various applications such as document classification, sentiment analysis and spam filtering. Numerous deep learning models have been proposed for text classification in the past few decades, and multilayer perceptron is one of the earliest architectures adopted to classify documents [133, 134]. The model typically has a single hidden layer with a number of units between 15 and 150. Text data is inherently sequential, as it is composed of a series of words and symbols arranged in a specific order. This property makes RNN and its variants particularly well-suited for processing and analyzing text data. In [135], RNN with two hidden layers, each with 6 units is used to classify news documents into eight classes. A study was conducted to investigate the variants of RNN i.e. LSTM and GRU for text classification [136]. The input to the model is a sequence of words of fixed length. The input sequence is also sliced into smaller subsequences of fixed length and passed to an independent model for parallelization. A convolutional layer can extract local features, allowing the model to leverage hierarchical temporal information in textual data. A hybrid model of convolutional and LSTM architecture is proposed for text classification [137]. Two parallel convolutional layers are used to extract features from word embeddings, followed by max-pooling layers to reduce the feature dimensions. The reduced features are then concatenated and passed to LSTM for prediction.
@@
Although CNN and RNN provide excellent results on text classification tasks, the models lack the ability to attend to specific words based on their importance and context. To address this limitation, attention mechanism is incorporated into the model to focus on the important features, enhancing the text classification accuracy. In [138], two attention modules are introduced to capture the contextual information of the feature sequence extracted by bi-directional LSTM. The first attention module attends the sequence in forward direction while the backward sequence is attended by the second attention module. The convolutional layers are used before the bi-directional LSTM to extract features from the word embedding. The attention modules require sequential processing using RNN-based architecture such as LSTM and GRU which may lead to information loss and distorted representations, particularly in long sequence. Furthermore, the attention modules focus on inter-sequence relationships between the input sequence and the target, ignoring the intra-sequence relationships or the dependencies between the words. In [139], the deep learning model is integrated with self-attention to capture the intra-sequence relationships between the features in the sequence. A multilayer of bi-directional LSTMs is utilized to extract feature sequence from the word embedding before the self-attention module attends the feature sequence to compute the attention weights. To further improve the overall performance, a multichannel features consisting of three input pipelines is introduced [140]. Each pipeline concatenates the word vector with a feature vector derived from the input sequence such as the word position, part-of-speech and word dependency parsing. The input pipeline is connected to bi-directional LSTM, followed by a self-attention module to learn the dependencies between the features in the sequence.
@@
The transformer is a deep learning architecture that transforms sequential data using self-attention mechanisms, allowing long-range dependencies and complex patterns to be captured. The architecture is the basis of various advanced deep learning models and the Bi-directional Encoder Representations from Transformers popularly known as BERT is one of the examples that leverage transformer for pre-training on large scale textual data [141]. BERT is a bi-directional transformer encoder that is designed for various NLP tasks, capable of capturing the contextual information from both preceding and succeeding words in the input sequence. Several improvements have been made to BERT to enhance its overall performance such as ALBERT [142], RoBERTa [143] and DeBERTa [144]. The improvements are centered around refining the pre-training approaches such as dynamic masking of the training instances, training with a block of sentences and representing each input word using two vectors, both content and position of the word. Most of the recent works leverage BERT and its variants to capture effective feature representation of the input sequence. In [145], BERT and its variants are leveraged to capture the long-range dependencies of the input tokens. The features are then passed to a layer normalization and a linear fully-connected layer with dropout for classification. Similar work is reported in [146] whereby BERT is used to extract the features and the features are then passed to a hybrid of convolutional and recurrent neural networks. The traditional machine learning algorithms have been used to classify the features extracted by BERT [147]. The study shows machine learning algorithms can effectively leverage the rich contextual features extracted by BERT for downstream classification tasks.
@@
In text classification, the text labels can help in capturing the words relevant to the classification. The label-embedding attentive model is one of the earliest attempts to joint learn the label and word embeddings in the same latent space and measure the compatibility between labels and words using cosine similarity [148]. The joint embedding allows the model to capture more effective text representations, increasing the overall performance of the model. LANTRN is a deep learning model that leverages label embedding extracted by BERT and entity information e.g. person name and organization name for text classification [149]. The entity recognition module is based on bi-directional LSTM and conditional random field layers to calculate the probability of each word in each entity label. The model introduces a label embedding bi-directional attention to learn the attention weights of token-label and sequence-label pairs. Furthermore, a transformer is introduced to learn local short-term dependencies of multiple short text sequences and long-term dependencies of the input sequence. Aspect refers to a specific attribute of an entity within the text and incorporating this information enhances the model’s understanding of the nuances of the text. BERT-MSL is a multi-semantic deep learning model with aspect-aware enhancement and four input pipelines: left sequence, right sequence, global sequence and aspect target [150]. The aspect-aware enhancement module takes the features extracted by BERT, and performs average pooling followed by a linear transform. Then the output is concatenated with the outputs produced by the local and global semantic learning modules. The concatenated features are then jointly attended by a multi-head attention for text classification. Figure 23 shows the architecture of BERT-MSL.

### v6

@@ 4.1.3 Image Segmentation
Image segmentation is another important task in which deep learning has a significant impact. One of the earliest deep learning models for image segmentation is the fully convolutional network [131]. A fully convolutional network consists of only convolutional layers which accepts an input of an arbitrary size and produce the predicted segmentation map of the same size. The authors adopted the AlexNet, VGG16 and GoogleNet, replace their fully connected layers with convolutional layers and append a 1×1 convolutional layer, followed by bilinear up-sampling to match the size of the input. The model was considered a significant milestone in image segmentation, demonstrating the feasibility of deep learning for semantic segmentation trained in end-to-end manner. Deconvolution network is another popular deep learning model for semantic segmentation [132]. The model architecture consists of two parts: encoder and decoder. The encoder takes an input image and uses the convolutional layers to generate the feature maps. The feature maps are fed to the decoder composed of un-sampling and deconvolutional layers to predict the segmentation map. SegNet is another encoder-decoder model for semantic segmentation [133]. The encoder is a sequence of convolutional (with ReLU) and maximum pooling blocks which is analogous to a convolutional neural network. The decoder is composed of up-sampling layers which up-samples the inputs using the memorized pooled indices generated in the encoder phase, and convolutional layers without non-linearity. The encoder progressively reduces the resolution of the input data while extracting abstract features through a series of convolutional and pooling layers. This process causes the loss of fine-grained information, degrading the overall performance of segmentation. LinkNet mitigates this limitation by passing the feature maps at several stages generated by the encoder to the decoder, hence reducing information loss [134]. The model architecture of LinkNet is similar to SegNet, but utilizes ResNet as the encoder.
@@
While Faster R-CNN is a significant approach in object detection task, it has been extended to perform instance segmentation task. One such extension is Mask R-CNN which is based on Faster R-CNN, introduces an additional branch for predicting the segmentation mask [135]. Similar to Faster R-CNN, Mask R-CNN utilizes the RPN to generate region proposals and then the region of interest alignment is applied to extract more accurate features from the proposed regions. Mask R-CNN does not leverage the multiscale feature representation which may degrade the overall performance of segmentation. To overcome this limitation, Path Aggregation Network (PANet) incorporates the FPN and introduces a bottom-up pathway to facilitate the propagation of the low-level information [136]. The pathway takes the feature maps of the previous stage as input and performs 3×3 convolution with stride 2 to reduce the spatial size of the feature maps. The generated feature maps are then fused with the feature maps from the FPN through the lateral connection. The model adopts the three branches as in Mask R-CNN. MaskLab is an instance segmentation model based on Faster R-CNN, consisting of object detection, segmentation, and instance (object) center direction prediction branches [137]. The direction prediction provides useful information to distinguish instances of the same semantic label, allowing the model to further refine the instance segmentation results.
@@
The convolutional layers use local receptive fields to process input data which can be effective for exploiting spatial patterns and hierarchical features but may find it difficult to capture global relationships across the entire image. The ViT has been leveraged to mitigate this issue in semantic segmentation [140]. Specifically, the input image is divided into patches and treated as input to the transformer to capture the long-range relationship between the patches, significantly improving the prediction of the segmentation map. Global context ViT aims to address the lack of ViT’s ability to leverage local spatial features [141]. As shown in Figure 24, the transformer consists of local and global self-attention modules. The role of global self-attention is to capture the global contextual information from different image regions while the short-range information is captured by the local self-attention. Multiscale feature representation is crucial for accurate semantic segmentation. However, the transformer often combines the features without considering their appropriate (optimal) scales, thus affecting the segmentation accuracy. Transformer scale gate is a module proposed to address the issue of selecting an appropriate scale based on the correlation between patch-query pairs [142]. The transformer takes attention (correlation) maps as input and calculates the weights of the multiscale features for each image patch, allowing the model to adaptively choose the optimal scale for each patch. Table 4 lists the summary of state-of-the-art image segmentation.
@@
Table 4: Summary of state-of-the-art image segmentation.
Ref. Description Results (Datasets and Metrics)
[138], 2023 The study proposed a multistage context refinement network with a context attention refinement module that improves feature representation by capturing local and global context, fusing information, and removing redundancies, with attention integrated into the encoder-decoder skip connections PASCAL VOC2012
mIoU: 79.05
ADE20K MI
mIoU: 41.25
Cityscapes
mIoU: 79.42
[139], 2022 This study introduced a covariance attention to model the local and global dependency for the feature maps by formulating a covariance matrix, providing complementary information to improve segmentation performance Cityscapes
mIoU: 82.9
Pascal-Context
mIoU: 51.7
ADE20K
mIoU: 43.78
[140], 2021 This study introduced a transformer for semantic segmentation Cityscapes
mIoU: 81.3
Pascal-Context
mIoU: 59.0
ADE20K
mIoU: 53.63
[141], 2023 This study proposed an improved vision transformer architecture that employes alternating global and local self-attentions to effectively capture both local and global spatial information ADE20K
mIoU: 49.2
[142], 2023 This study introduced TSG module that uses attention maps to calculate multiscale feature weights for each image patch, enabling the model to adaptively select the optimal scale based on patch-query correlation Cityscapes
mIoU: 83.6
Pascal-Context
mIoU: 64.9
ADE20K
mIoU: 56.93
@@ 4.1.4 Image Generation
Image generation refers to the process of creating images based on input texts. Generally, the task can be divided into three stages. The first stage is extracting features from the input text, followed by generating the image and finally controlling the image generation process to ensure the output meets specific criteria and constraints. This section focuses on the progress made in the development of deep learning models of the second stage (image generation) since it directly impacts the quality of the generated images. Variational autoencoder is one of the earliest deep learning models that is capable of generating images [84]. Variational autoencoder learns to generate data by capturing the underlying (Gaussian) distribution of the training data. During the generation process, the distribution parameters are sampled and passed to the decoder to generate the output image. Although the generated images are blurry and unsatisfactory, it has shown a lot of potential in image generation tasks. The introduction of GAN significantly improved the quality of generated images. GAN consists of two connected neural networks, a generator and a discriminator that are trained simultaneously in a competitive manner [85]. The generator learns to generate realistic images to fool the discriminator, while the discriminator learns to distinguish between fake and real images. The generated images are less blurry and more realistic. Several enhanced models have been proposed to improve its usability and overall performance such as conditional GAN [86] which allows us to tell what image to be generated, and the deep convolutional GAN (DCGAN) [143] which provides a more stable structure for image generation. DCGAN is the basis of many subsequent improvements in GANs.
@@
Although multistage generators improve image generation performance by leveraging multiscale representation, the generated images may contain fuzzy shapes with coarse features. DF-GAN replaces the multistage generators with a single-stage deep generator featuring residual connections and trained with hinge loss [151]. Furthermore, DF-GAN introduces a regularization strategy on the discriminator that applies a gradient penalty on real images with matching text, allowing the model to generate more text-matching images. DMF-GAN an improved DF-GAN, incorporates three novel components designed to leverage semantic coherence between the input text and the generated image [152]. The first component is the recurrent semantic fusion module, which models long-range dependencies between the fusion blocks. The second component is the multi-head attention module, which is placed towards the end of the generator to leverage the word features, forcing the generator to generate images conditioned on the relevant words. The last component is the word-level discriminator, which provides fine-grained feedback to the generator, facilitating the learning process and improving the overall quality of the generated images. Figure 26 shows the architecture of DMF-GAN. The process of image generation involves feeding a noise vector to the generator at the very beginning of the network. However, as the generator goes deeper, the noise effect may be diminished, affecting the diversity of the image generation results. To mitigate this issue, DE-GAN incorporates a dual injection module into the single-stage generator [153]. The dual injection module consists of two text fusion layers followed by a noise broadcast operation. The text fusion layer takes the sentence embedding and fuses it with the input feature map using the fully connected layer. Then noise is injected into the output feature map to retain the randomness in the generation process, improving diversity and generalization of the model. Table 5 lists the summary of state-of-the-art image generation.
@@
Table 5: Summary of state-of-the-art image generation.
Ref. Description Results (Datasets and Metrics)
[151], 2022 This study introduced a one-stage GAN architecture, employing a Target-Aware Discriminator to enhance text-image semantic consistency and utilizes deep text-image fusion blocks for effective feature integration, resulting in improved image authenticity and relevance CUB
IS: 5.10, FID: 14.81
COCO
FID: 19.32
[152], 2024 This study proposed a deep multimodal fusion generative adversarial networks that allow effective semantic interactions for fusing text information into the image synthesis process CUB
IS: 5.42, FID: 13.21
COCO
IS: 36.72, FID: 15.83
[153], 2024 This study proposed a deep learning-based method which utilizes BERT to extract sentiment information and analyze sentiment at the sentence and word level with SVM for sentiment classification CUB
IS: 4.86, FID: 18.94
COCO
IS: 18.33, FID: 28.79
@@ 4.2 Natural Language Processing
Natural language processing (NLP) refers to the field of AI that concerns with enabling computers to process, analyze and interpret human languages to extract useful information. Some of the common tasks in NLP are machine translation, text classification and text generation. Deep learning has been widely applied to solve real-world NLP problems. This section presents the recent advancements in deep learning models that have been designed for NLP over the past few years.
@@ 4.2.1 Text Classification
Text classification, known as text categorization, is a task that involves assigning predefined categories or labels to a piece of text based on its content. The task is commonly used in various applications such as document classification, sentiment analysis and spam filtering. Numerous deep learning models have been proposed for text classification in the past few decades, and multilayer perceptron is one of the earliest architectures adopted to classify documents [154, 155]. The model typically has a single hidden layer with a number of units between 15 and 150. Text data is inherently sequential, as it is composed of a series of words and symbols arranged in a specific order. This property makes RNN and its variants particularly well-suited for processing and analyzing text data. In [156], RNN with two hidden layers, each with 6 units, is used to classify news documents into eight classes. A study was conducted to investigate the variants of RNN such as LSTM and GRU for text classification [157]. The input to the model is a sequence of words of fixed length. The input sequence is also sliced into smaller subsequences of fixed length and passed to an independent model for parallelization. A convolutional layer can extract local features, allowing the model to leverage hierarchical temporal information in textual data. A hybrid model of convolutional and LSTM architecture is proposed for text classification [158]. Two parallel convolutional layers are used to extract features from word embeddings, followed by maximum pooling layers to reduce the feature dimensions. The reduced features are then concatenated and passed to LSTM for prediction.
@@
Although CNN and RNN provide excellent results on text classification tasks, the models lack the ability to attend to specific words based on their importance and context. To address this limitation, an attention mechanism is incorporated into the model to focus on the important features, enhancing the text classification accuracy. In [159], two attention modules are introduced to capture the contextual information of the feature sequence extracted by bidirectional LSTM. The first attention module attends the sequence in forward direction, while the backward sequence is attended by the second attention module. The convolutional layers are used before the bidirectional LSTM to extract features from the word embedding. The attention modules require sequential processing using RNN-based architecture such as LSTM and GRU, which may lead to information loss and distorted representations, particularly in long sequence. Furthermore, the attention modules focus on inter-sequence relationships between the input sequence and the target, ignoring the intra-sequence relationships or the dependencies between the words. In [160], the deep learning model is integrated with self-attention to capture the intra-sequence relationships between the features in the sequence. A multilayer of bidirectional LSTMs is utilized to extract the feature sequence from the word embedding before the self-attention module attends the feature sequence to compute the attention weights. To further improve the overall performance, a multichannel features consisting of three input pipelines is introduced [161]. Each pipeline concatenates the word vector with a feature vector derived from the input sequence such as the word position, part-of-speech and word dependency parsing. The input pipeline is connected to bidirectional LSTM, followed by a self-attention module to learn the dependencies between the features in the sequence.
@@
The transformer is a deep learning architecture that transforms sequential data using self-attention mechanisms, allowing long-range dependencies and complex patterns to be captured. The architecture is the basis of various advanced deep learning models and the Bidirectional Encoder Representations from Transformers popularly known as BERT is one of the examples that leverage transformer for pre-training on large scale textual data [162]. BERT is a bidirectional transformer encoder that is designed for various NLP tasks, capable of capturing the contextual information from both preceding and succeeding words in the input sequence. Several improvements have been made to BERT to enhance its overall performance, such as ALBERT [163], RoBERTa [164] and DeBERTa [165]. The improvements are centered around refining the pre-training approaches such as dynamic masking of the training instances, training with a block of sentences and representing each input word using two vectors, both content and position of the word. Most of the recent works leverage BERT and its variants to capture effective feature representation of the input sequence. In [166], BERT and its variants are leveraged to capture the long-range dependencies of the input tokens. The features are then passed to a layer normalization and a linear fully connected layer with dropout for classification. Similar work is reported in [167] whereby BERT is used to extract the features and the features are then passed to a hybrid of convolutional and recurrent neural networks. The traditional machine learning algorithms have been used to classify the features extracted by BERT [168]. The study shows machine learning algorithms can effectively leverage the rich contextual features extracted by BERT for downstream classification tasks.
@@
In text classification, the text labels can help in capturing the words relevant to the classification. The label-embedding attentive model is one of the earliest attempts to joint learn the label and word embeddings in the same latent space and measure the compatibility between labels and words using cosine similarity [169]. The joint embedding allows the model to capture more effective text representations, increasing the overall performance of the model. LANTRN is a deep learning model that leverages label embedding extracted by BERT and entity information e.g. person name and organization name for text classification [170]. The entity recognition module is based on bidirectional LSTM and conditional random field layers to calculate the probability of each word in each entity label. The model introduces a label embedding bidirectional attention to learn the attention weights of token-label and sequence-label pairs. Furthermore, a transformer consisting of RNN and a multi-head self-attention mechanism is introduced to learn local short-term dependencies of multiple short text sequences and long-term dependencies of the input sequence. Aspect refers to a specific attribute of an entity within the text, and incorporating this information enhances the model’s understanding of the nuances of the text. BERT-MSL is a multi-semantic deep learning model with aspect-aware enhancement and four input pipelines: left sequence, right sequence, global sequence and aspect target [171]. The aspect-aware enhancement module takes the features extracted by BERT, and performs average pooling followed by a linear transform. Then the output is concatenated with the outputs produced by the local and global semantic learning modules. The concatenated features are then jointly attended by a multi-head attention for text classification. Figure 27 shows the architecture of BERT-MSL. Table 6 lists the summary of state-of-the-art text classification.
@@
Table 6: Summary of state-of-the-art text classification.
Ref. Description Results (Datasets and Metrics)
[166], 2022 The study proposed an enhanced RoBERTa by incorporating layer normalization and dropout to enable smoother gradients, faster training and reduced overfitting, thus, improving model generalization NEISS
Accuracy: 0.98
[167], 2024 The study employed IndoBert to extract features from text data which were then passed to a hybrid deep learning model for sentiment analysis Shopee
Accuracy: 0.8612 (CNN-GRU)
Tokopedia
Accuracy: 0.8768 (LSTM-CNN)
Lazada
Accuracy: 0.8710 (LSTM-CNN)
[168], 2023 This study proposed a deep learning-based method which utilizes BERT to extract sentiment information and analyze sentiment at the sentence and word level with SVM for sentiment classification NLPCC-SCDL
Accuracy: 95.12%
[170], 2023 This study proposed LANRTN model for text classification, integrating an R-Transformer with label embedding, attention mechanisms, and an entity recognition model to capture both global dependencies and label-aware contextual information Reuters Corpus Volume I
Micro F1-score: 0.893, Arxiv Academic Paper Dataset
Micro F1-score: 0.718
[171], 2023 The paper proposed a BERT-MSL model for aspect-based sentiment analysis, integrating multiple semantic learning modules and merging them using multi-head self-attention and linear transformation layers for enhanced classification 14Lap
Accuracy: 82.24%, Macro-average F1-score: 78.98%
14Rest
Accuracy: 89.11%, F1-score: 84.07%
15Rest
Accuracy: 88.53%, Macro-average F1-score: 71.41%
16Rest
Accuracy: 93.09%, Macro-average F1-score: 81.24%
Twitter
Accuracy: 74.89%, Macro-average F1-score: 73.43%